OpenStudy (unklerhaukus):

By explicit construction of the matrices in question, show that any matrix T can be written as the sum of matrices: a) symmetric matrix S and antisymmetric matrix A b) real matrix R and imaginary matrix M c) hermitian matrix H and skew-hermitian matrix K

OpenStudy (jamesj):

are you stuck on this, really? Which bit? part b is trivial yes?

OpenStudy (unklerhaukus):

i have the answer i just dont find it particularly illuminating

OpenStudy (jamesj):

Hmm ... for part a it's A = (1/2) (A + A^t) + (1/2) (A - A^t) I think that's quite intuitive.

OpenStudy (jamesj):

In fact, all three parts have the 'same' answer if you replace transpose for conjugate for complex transpose

OpenStudy (jamesj):

It's analogous also to writing a real function as the sum of an even and an odd function.

OpenStudy (unklerhaukus):

the answer for a) i have is \[S=\frac{1}{2}(T+\tilde{T}); \quad A=\frac{1}{2}(T-\tilde{T})\]

OpenStudy (jamesj):

right, where ~ means transpose

OpenStudy (unklerhaukus):

yeah

OpenStudy (unklerhaukus):

ok well is there a step before this that i can not see for some reason

OpenStudy (jamesj):

Um, I arrived at it by remembering that A + A^t is always symmetric. Then I just had to scale the thing appropriately and find it's complement. Then I remembered the pattern works everywhere

OpenStudy (jamesj):

*its

OpenStudy (unklerhaukus):

So i guess my question should be why is \[A +\tilde {A} \]aka\[A+A^T\] necessarily symmetric

OpenStudy (jamesj):

Because (A^t)^t = A, hence .... see it now?

OpenStudy (unklerhaukus):

\[A+A^T=(A^T+A)^T\]\[A^T+A = (A^T+A)^T\]

OpenStudy (unklerhaukus):

the bottom right corner i mean to have\[(A+A^T)^T\]

OpenStudy (unklerhaukus):

and the point is transposes do noting to the sum.

OpenStudy (unklerhaukus):

have i got it?

OpenStudy (jamesj):

Yes. A little more abstractly, the operation of taking the transpose is order 2. This is why the same method will also work with the other operations.

OpenStudy (jamesj):

Remember when we were kids in primary school we'd make painting of butterflies

OpenStudy (jamesj):

We'd put a vertical fold in the paper and put some paint on both sides. Then we'd fold the two sides together and open it up again. The analogy isn't perfect, but that's sort of what's going on here.

OpenStudy (jamesj):

On that colourful note, I'm out of here.

OpenStudy (unklerhaukus):

thanks i think it makes more sense to me now

OpenStudy (unklerhaukus):

this is a bit like odd and even functions isnt it